EXERCISE 2 


The following three exercises do belong together. Assume that we have five assets. The first one has expected return μ1 = 20% and standard deviation of return equal to σ1 = 10%. The second has expected return μ2 = 40% and standard deviation of return equal to σ2 =20%. 


Assume that the third asset has expected return μ3 = 10%. What is the range of the standard deviation σ3 of the third asset so that the three assets form an efficient set?


Select one:


a. 


The standard deviation of the third asset needs to be below 10%.


b. 


The range is the empty set, as it is not possible that all three assets are efficient in this case.


c. 


The standard deviation of the third asset needs to be above 10%.


d. 


The standard deviation of the third asset needs to be below 20%.


e. 


 The standard deviation of the third asset needs to be above 20%.


f. 


The standard deviation of the third asset needs to be between 10% and 20%.



1
Expert's answer
2021-05-25T18:37:54-0400

As For asset 1 CV = 20 / 10 = 2

For 2 Asset 2 CV = 40/ 20 = 2

For asset 3 CV should be at 2 S.D. = 10 / 2 = 5%

Therefore, the standard deviation of the third asset needs to be below 10%.